# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

### v2

@@ 4.2. Inlier and Outlier Weight Quantization
Each row to be quantized is first divided into multiple non overlapping contiguous macro-blocks (MaBs) of size $B_{M}=128$. All inliers are quantized within a MaB and share the scale factor. Each MaB is then subdivided into multiple non-overlapping contiguous micro-blocks ($\mu$Bs) of size $B_{\mu}=8$ with sixteen $\mu$Bs forming a MaB. The outliers present in each $\mu$B shares same scale(s). As depicted in Figure 3(a), tab 1, the quantization process begins by first identifying inliers and outliers in each MaB by using the $3\sigma$ rule. A shared 8-bit power-of-two scale factor ($2^{I_{sf}}$), following Equation 1 is calculated for all inliers in a MaB and the inliers are quantized to 2-bit or 4-bit integers resulting in MX-INT-(2/4)128 quantization. Interestingly, we observe that the inlier scale factor in each MaB is always a negative power of two for all FMs under consideration. We leverage this observation to reduce outlier magnitude, by multiplying all outlier values in a MaB with the inlier scale factor ($2^{I_{sf}}$) (can also be perceived as division by $2^{-I_{sf}}$, for conformity with Equation 2). This pre-processing helps make outlier quantization easier, by pre-reducing its maximum dynamic range.
Unlike inliers, outliers are quantized per $\mu$B, to reduce quantization error due to shared scaling over a larger group size (see section 6). After identifying outliers present in a $\mu$B, we compute a shared 8-bit MXScale that is calculated by concatenating the level-1 power-of-two scale factor ($2^{O^{l_{1}}_{sf}}$) and level-2 microExponent ($\mu X$). The level-1 scale factor is calculated by following Equation 1 to obtain 7 or 5-bit MSBs of MXScale depending on size of $\mu X$ being 1 or 3-bit of the LSBs–corresponds to exponent size of the FP format (depicted in tab 2 in Figure 3(a)). The outliers in a $\mu$B are scaled by ($2^{O^{l_{1}}_{sf}}$), following Equation 2 and then quantized to either e1m2/e3m4 FP-format for $b$ = 4 or 8-bit, respectively. Post quantization of outliers, the level-two scale factor or the $\mu X$ is obtained by extracting the common exponent among all outliers in a $\mu$B. This process results in a MX-FP-$b$8,8 outlier quantization. The final outlier scale factor is $2^{O_{sf}}$ where, $O_{sf}$ is expressed as ${O_{sf}}={O^{l_{1}}_{sf}+\mu X-I_{sf}}$. The term $I_{sf}$ in the final outlier scale factor accounts for multiplication by $2^{I_{sf}}$ (or division by the inverse) during outlier pre-processing.
@@
5             # Divide each row into non-overlapping Macro-Blocks
@@
7                   # Separate inlier and outlier in each Macro-Block
@@
9                  # Quantize Inliers to lower precision
@@ 10                  $\mathbb{Q}^{in},I_{sf}$ = InlierQuantization($\mathbb{W}^{in}$)
12                         # Count Number of Outliers in a Micro-Block
@@
14                        # Initialize Inlier Index List
@@
16                         # Identify n least Important Inlier Position
@@ 19                              $p=argmin_{p\in\mathbb{W}^{in}_{\mu B}}w_{p}^{2}/[\mathbb{H}^{-1}]_{pp}$
20                              # Prune least important Inliers
@@ 21                              $w_{p}\leftarrow 0$
22                              # Update M with the location of $w_{p}$
@@ 23                              $M\leftarrow M+\{p\}$
26                        # Quantize Outliers to higher precision
@@ 27                        $\mathbb{Q}^{out},O_{sf}$ = OutlierQuant($\mathbb{W}^{out}_{\mu B},I_{sf}$)
28                        # Distribute LSB Outlier Bits to Sparse Inlier Indices
@@ 36            $\mathbb{E}_{(j-i),:}=(\mathbb{W}_{j,:}-\mathbb{Q}_{j,:})/[\mathbb{H}^{-1}]_{jj}$
37            # Update weights in rB to compensate quantization error
@@
40      # Update remaining weights after a row block is quantized
@@ 4.3. Outlier Value Encoding through N:M Structured Pruning
Let us assume $n$ outliers are present in a $\mu$B quantized to MX-FP-44448,8. After sharing 1-bit $\mu X$ across the $n$ outliers, each outlier has 1 sign ($s$) and 2 mantissa bits ($m_{1}m_{0}$). The inliers are 2-bit 2’s complement MX-INT and has 1-bit sign and 1-bit magnitude. Since the outlier bits have two mantissa bits with one sign, to ensure symmetric outlier distribution, we duplicate their sign bit and assign each sign to a mantissa creating and partitioning into two halves Upper, Lower of size 2-bit each mimicking inlier MX-INT structure i.e., $\{sm_{1},sm_{0}\}$ (see Figure 3(a), tab 2). While we demonstrate this for a specific example, the process can be generalized to other inlier and outlier formats bit-widths, wherein the remaining outlier bits need to be reorganized into Upper and Lower distributable halves each of size $bb$ (per element bit-budget), with $bb$ representing the operational bit-width of each PE.
To ensure a fixed $bb$ and data-type of a layer for aligned memory access and simple PE design, we distribute the outlier LSBs (Lower half) to the least important inlier locations that are pruned within a $\mu$B. For $n$ outliers in a $\mu$B we identify $n$ least important inlier locations to prune via the Hessian information (see Algorithm 1), forming a ($B_{\mu}$-$n$):$B_{\mu}$ structured pruning pattern (Jeong et al., 2023; Sun et al., 2021) i.e., ($B_{\mu}$-$n$) non-zero values exist per $\mu$B after pruning. To keep track of the corresponding halves of each of the outliers in a $\mu$B, we maintain a per-$\mu$B permutation list (Note: if there are no outliers in a $\mu$B, a permutation list is not stored). The permutation list for each $\mu$B that has outliers is made up of $\frac{B_{\mu}}{2}$ elements (maximum number of outliers supported per $\mu B$) with each element storing the locations of the Upper and Lower halves of outliers in a 6-bit format $\{\texttt{Upper}_{loc},\texttt{Lower}_{loc}\}$ (see Figure 3(a) tab 3). Note that, in case of outlier count $>\frac{B_{\mu}}{2}$, the mB size should be chosen to be higher to prevent pruning of outliers and causing higher accuracy degradation. While such a situation does not arise in any of the models that we evaluate with $B_{\mu}=8$, this feature only serves as a demonstration of the flexibility of MicroScopiQ for future model architectures.
Figure 4. Comparison of memory organization between OliVe and MicroScopiQ.
4.4. Effective Bit Width Calculation
@@
5. MicroScopiQ Architecture
Figure 5. Integration of MicroScopiQ into a WS systolic array. The numbering in blue denotes the computation pipeline of MicroScopiQ and (see section 5.2)
5.1. Memory Organization
In Figure 4(b), we show how a typical weight layer is organized in memory (DRAM/HBM) for acceleration by MicroScopiQ inline with other multi-precision accelerators (Ramachandran et al., 2024b; Dai et al., 2021; Keller et al., 2023; Guo et al., 2023) as shown in Figure 4(a) for OliVe. It is comprised of two sections, quantized weights and metadata. For MicroScopiQ-W4, all weight elements occupy 4-bits and assuming 1-byte as the smallest addressable memory (Guo et al., 2023), two adjacent weight elements occupy a memory word and aligns 8-bit memory accesses with high efficiency. The metadata section consists of 8-bit inlier scale factors which are shared per MaB and for a $\mu B$ of size 8, there are 16 $\mu Bs$. A 1-bit identifier per $\mu B$ is used to notify the controller whether a particular $\mu B$ has outliers and if the corresponding metadata (8-bit MXScale and 24-bit permutation list) needs to be fetched. The controller calculates how many $\mu B^{\prime}s$ metadata are stored in memory and burst accesses them accordingly and stores it in the instruction buffer. All the metadata is also memory word aligned, in line with the definition proposed in (Guo et al., 2023). Following the tiling, banking scheme of popular accelerators, (Chen et al., 2019; Keller et al., 2023; Muñoz-Martínez et al., 2023; Dai et al., 2021) we employ dedicated interleaved SRAM banks for each row for weights and activations (circularly rotated across banks) (Peemen et al., 2013) and a separate bank for the metadata, enabling it to be loaded in parallel. In WS systolic arrays, the weight matrix W is spatially laid out onto the systolic arrays; thus, W must be partitioned into tiles of R × C to match the array dimensions, where R and C denote the number of array rows and columns, respectively. Because the first dimension of W must match the second dimension of the activation matrix X in a matrix multiplication, X’s second dimension is also required to be partitioned with a size of R.
@@ 5.1. Memory Organization
Figure 6. Tiling and scheduling of a GEMM operation on a $2\times 2$ MicroScopiQ accelerator.
5.2. Computation Flow
In Figure 5, we depict the MicroScopiQ accelerator architecture, integrated into a standard weight stationary systolic array. The accelerator is optimized for throughput and all compute units (PE array, ReCoN etc.) are internally pipelined with interleaved pipeline stages. Through the numbered steps in Figure 5, we explain the computational flow of the MicroScopiQ accelerator. Assume a quantized FM with potential outliers conditioned and distributed within a $\mu$B and corresponding metadata calculated offline. 1 One $\mu$B or multiple $\mu$Bs of weights are mapped to each PE row as typically $B_{\mu}\leq$#\## of columns in PE array (Jouppi et al., 2017)  [^5] To account for ##\## of columns $<$ $B_{\mu}$, $\mu$Bs are tiled and mapped to the PE array iteratively. The merging of outlier halves is then conducted separately from the partial sum calculation through another computation cycle. Each PE in a row either receives a weight at high-precision (e.g. 4-bit) or multiple weights with low-precision (e.g. two 2-bits). 2 PE row 0 receives quantized input activation (iAct) from left and input accumulation (iAcc) from top. 3 The PEs in row 0 perform INT multiplication of the stationary weights with the iActs, irrespective of outlier/inlier weights. During accumulation, assuming the $\mu$Bs mapped to PE row 0 do not contain outliers, the PEs accumulate the computed multiplication result with iAcc and direct the partial sums to PE row 1. 4 PE row 1 receives partial sums from top and similarly performs INT multiplication between weights and iAct. In parallel, the output scale factors can be calculated as described in section 5.6. 5 Assuming the mapped $\mu$Bs to PE row 1 have atleast one outlier, the controller directs all PE outputs to ReCoN. The PE with outliers offloads the accumulation to ReCoN by sending a concatenation of the multiplication result (Res) and iAcc at its output. This is needed because the outlier will have its two halves distributed in different locations and possibly mapped to different PEs. While the PEs in row 1 with inlier weights can accumulate the partial sum output, the PEs with the Upper and Lower outlier halves cannot compute the outlier partial sum output. 6 After receiving outputs from row 1, ReCoN reorders and calculates the FP-outlier partial sum, leaving the inlier partial sums untouched. 7 The controller signals PE row 2 to expect input partial sums from ReCoN and not from the previous row. Similarly computations proceed and finally producing output activations (oActs). 8 oActs are post-processed (scaled and quantized). The post-processing of oActs does not require all the oActs to be computed. The post processing operations can be conducted as and when the oActs are generated and overlapped with the rest of the computation so as to hide computation and memory access latency overhead. In situations involving back-to-back GEMMs, we follow the standard practice adopted by SoTA accelerators (Keller et al., 2023; Shao et al., 2019; Chen et al., 2019; Ramachandran et al., 2024b), wherein the oActs are either directly routed to the iAct buffer or staged in L2 memory (based on the buffer capacity and oAct size) so as to eventually load into the iAct buffer for the next computation. (shown by the green dotted line in Figure 5).
In Figure 6, through a small-scale example on a $2\times 2$ MicroScopiQ array, we demonstrate how iActs and weights are tiled and scheduled for processing across space and time. In Figure 6, row 0 of mapped weights has only inliers and row 1 of mapped weights has a single outlier with its upper and lower halves occupying two spatial locations.
@@ 5.3. On-chip Storage and Control
MicroScopiQ Controller. The controller generates appropriate control and signals to all units. It is functionally very similar to standard systolic array controllers (Sharma et al., 2018), however, with added functionality to support specific features of ReCoN, multi-precision PEs, post-processing unit. Furthermore, unlike traditional controllers, the MicroScopiQ controller exerts fine-grained control using handshaking signals on the streaming of iActs and iAccs into the PE array to account for the increased pipeline depth through ReCoN for PE rows with outliers.
5.4. Multi-Precision INT-PE Array
Multi-Precision. Existing multi-precision accelerators (Guo et al., 2023; Sharma et al., 2018; Guo et al., 2022) typically follow a bottom-up approach, employing low-precision PEs and grouping neighboring PEs to support higher precision. This sacrifices throughput for multi-precision support, as multiple PE columns are required to perform a single MAC operation. This reduces parallelism and increases latency. We adopt a different strategy via the MODE signal for multi-precision support by mapping multiple weights that share the same rightbound iAct to the same PE at lower precision (2-bits) or a single weight at higher precision (4-bits). The two 2-bit weights mapped to the same PE in MicroScopiQ are the weights that would have been mapped to different columns within the same row in existing multi-precision accelerators. At high-precision operation we utilize available parallelism from the PE array while at lower-precision operation we increase computation throughput by facilitating the parallel evaluation of multiple partial sums along each column.
Multiplication Stage. Inspired by (Liu et al., 2023; Umuroglu et al., 2018) we present a partial product multiplier-tree architecture for multi-precision INT multiplication. The weights and iAct are partitioned across the four $4$-bit $\times$ $2$-bit multipliers to calculate partial sums ($P_{00},P_{01},P_{10},P_{11}$) as described in Figure 5. Based on the MODE signal the partial sums above are combined using a combination of adders and shifters to provide the result (Res) of different bit-precision weight with iAct as follows,

### v3

@@ 4.2. Inlier and Outlier Weight Quantization
In step 1, each row to be quantized is first divided into multiple non overlapping contiguous macro-blocks (MaBs) of size $B_{M}=128$. All inliers are quantized within a MaB and share the scale factor. Each MaB is then subdivided into multiple non-overlapping contiguous micro-blocks ($\mu$Bs) of size $B_{\mu}=8$ with sixteen $\mu$Bs forming a MaB. The outliers present in each $\mu$B shares same scale(s). As depicted in Figure 3(a), step 1, the quantization process begins by first identifying inliers and outliers in each MaB by using the $3\sigma$ rule. A shared 8-bit power-of-two scale factor ($2^{I_{sf}}$), following Equation 1 is calculated for all inliers in a MaB and the inliers are quantized to 2-bit or 4-bit, resulting in MX-INT-(2/4)128 quantization. Interestingly, we observe that the inlier scale factor in each MaB is always a negative power of two for all FMs under consideration. We leverage this observation to reduce outlier magnitude, by multiplying all outlier values in a MaB with the inlier scale factor ($2^{I_{sf}}$) (can also be perceived as division by $2^{-I_{sf}}$, for conformity with Equation 2). This pre-processing helps make outlier quantization easier, by pre-reducing its dynamic range.
Unlike inliers, outliers are quantized per $\mu$B, to reduce quantization error due to shared scaling over a larger group size (see section 6). After identifying outliers present in a $\mu$B, we compute a shared 8-bit MXScale that is calculated by concatenating the level-1 power-of-two scale factor ($2^{O^{l_{1}}_{sf}}$) and level-2 microExponent ($\mu X$). The level-1 scale factor is calculated by following Equation 1 to obtain 7 or 5-bit MSBs of MXScale depending on size of $\mu X$ being 1 or 3-bit of the LSBs–corresponds to exponent size of the FP format (depicted in step 2 in Figure 3(a)). The outliers in a $\mu$B are scaled by ($2^{O^{l_{1}}_{sf}}$), following Equation 2 and then quantized to either e1m2/e3m4 FP-format for $b$ = 4 or 8-bit, respectively. Post quantization of outliers, the level-two scale factor or the $\mu X$ is obtained by extracting the common exponent among all outliers in a $\mu$B. This process results in a MX-FP-$b$8,8 outlier quantization. The final outlier scale factor is $2^{O_{sf}}$ where, $O_{sf}$ is expressed as ${O_{sf}}={O^{l_{1}}_{sf}+\mu X-I_{sf}}$. The term $I_{sf}$ in the final outlier scale factor accounts for multiplication by $2^{I_{sf}}$ (or division by the inverse) during outlier pre-processing.
@@
5             # Step 1.0: Divide each row into non-overlapping Macro-Blocks
@@
7                   # Step 1.1: Separate inlier and outlier in each Macro-Block
@@
9                  # Step 1.2: Quantize Inliers to lower precision
@@ 10                  $\mathbb{Q}^{in},I_{sf}$ = InlierQuantization($\mathbb{W}^{in}$)
12                         # Step 2.0: Count Number of Outliers in a Micro-Block
@@
14                        # Step 2.1: Initialize Inlier Index List
@@
16                         # Step 2.2: Identify n least Important Inlier Position
@@ 19                              $p=argmin_{p\in\mathbb{W}^{in}_{\mu B}}w_{p}^{2}/[\mathbb{H}^{-1}]_{pp}$
20                              # Step 2.3: Prune least important Inliers
@@ 21                              $w_{p}\leftarrow 0$
22                              # Step 2.4: Update M with the location of $w_{p}$
@@ 23                              $M\leftarrow M+\{p\}$
26                        # Step 2.5: Quantize Outliers to higher precision
@@ 27                        $\mathbb{Q}^{out},O_{sf}$ = OutlierQuant($\mathbb{W}^{out}_{\mu B},I_{sf}$)
28                        # Step 3.0: Distribute LSB Outlier Bits to Sparse Inlier Indices
@@ 36            $\mathbb{E}_{(j-i),:}=(\mathbb{W}_{j,:}-\mathbb{Q}_{j,:})/[\mathbb{H}^{-1}]_{jj}$
37            # Step 3.2: Update weights in rB to compensate quantization error
@@
40      # Step 3.3: Update remaining weights after a row block is quantized
@@ 4.3. Outlier Value Encoding through N:M Structured Pruning
To formalize step 2, let us assume $n$ outliers are present in a $\mu$B quantized to MX-FP-44448,8. After sharing 1-bit $\mu X$ across the $n$ outliers, each outlier has 1 sign ($s$) and 2 mantissa bits ($m_{1}m_{0}$). The inliers are 2-bit 2’s complement MX-INT and has 1-bit sign and 1-bit magnitude. Since the outlier bits have two mantissa bits with one sign, to ensure symmetric outlier distribution, we duplicate their sign bit and assign each sign to a mantissa creating and partitioning into two halves Upper, Lower of size 2-bit each mimicking inlier MX-INT structure i.e., $\{sm_{1},sm_{0}\}$ (see Figure 3(a), step 2). While we demonstrate this for a specific example, the process can be generalized to other inlier and outlier formats bit-widths, wherein the remaining outlier bits need to be reorganized into Upper and Lower distributable halves each of size $bb$ (per element bit-budget), with $bb$ representing the operational bit-width of each PE.
In step 3, to ensure a fixed $bb$ and data-type of a layer for aligned memory access and simple PE design, we distribute the outlier LSBs (Lower half) to the least important inlier locations that are pruned within a $\mu$B. For $n$ outliers in a $\mu$B we identify $n$ least important inlier locations to prune via the Hessian information (see Algorithm 1), forming a ($B_{\mu}$-$n$):$B_{\mu}$ structured pruning pattern (Jeong et al., 2023; Sun et al., 2021) i.e., ($B_{\mu}$-$n$) non-zero values exist per $\mu$B after pruning. To keep track of the corresponding halves of each of the outliers in a $\mu$B, we maintain a per-$\mu$B permutation list (Note: if there are no outliers in a $\mu$B, a permutation list is not stored). The permutation list for each $\mu$B that has outliers is made up of $\frac{B_{\mu}}{2}$ elements (maximum number of outliers supported per $\mu B$) with each element storing the locations of the Upper and Lower halves of outliers in a 6-bit format $\{\texttt{Upper}_{loc},\texttt{Lower}_{loc}\}$ (see Figure 3(a) step 3). Note that, in case of outlier count $>\frac{B_{\mu}}{2}$, the mB size should be chosen to be higher to prevent pruning of outliers and causing higher accuracy degradation. While such a situation does not arise in any of the models that we evaluate with $B_{\mu}=8$, this feature only serves as a demonstration of the flexibility of MicroScopiQ for future models.
Figure 4. Integration of MicroScopiQ into a WS systolic array.
4.4. Effective Bit Width (EBW) Calculation
@@
5. MicroScopiQ Accelerator
@@
Figure 5. Example scheduling of a GEMM operation on a $2\times 2$ MicroScopiQ accelerator.
5.1. Architecture Overview
In Figure 4, we depict the MicroScopiQ accelerator architecture, integrated into a standard weight stationary systolic array. The accelerator is optimized for throughput and all compute units (PE array, ReCoN etc.) are internally pipelined with interleaved pipeline stages. Through the numbered steps in Figure 4, we explain the computational flow of the MicroScopiQ accelerator. And in Figure 5(a-c), we demonstrate the mapping and schedule for a sample $2\times 2$ GEMM operation with inlier and outlier weights, following the same computation flow.
Assume a quantized FM with outliers distributed within a $\mu$B and corresponding metadata calculated offline.
1
One $\mu$B or multiple $\mu$Bs of weights are mapped to each PE row as typically $B_{\mu}\leq$#\## of columns in PE array (Jouppi et al., 2017). Each PE in a row either receives a weight at high-precision (e.g. 4-bit) or multiple packed weights at low-precision (e.g. two 2-bits).
2
PE row 0 receives quantized input activation (iAct) from left and input accumulation (iAcc) from top. The PEs in row 0 perform multiplication of the stationary weights with the iActs, regardless of outlier/inlier weights. During accumulation, assuming the $\mu$Bs mapped to PE row 0 do not contain outliers (see Figure 5(c)), the PEs accumulate the computed multiplication result with iAcc and direct the partial sums to PE row 1.
3
PE row 1 receives partial sums from the top and similarly performs INT multiplication between weights and iAct. As shown in the scheduling figure, in parallel, the output scale factors can be calculated (section 5.5). 4 The mapped $\mu$Bs to PE row 1 have outliers, the controller directs all PE outputs to ReCoN. PE row 1 offloads the accumulation to ReCoN (shown by dotted lines in Figure 5(c)). This is needed because the outlier will have its two halves distributed in different locations and mapped to different PEs. While the PEs in row 0 with inlier weights can accumulate the partial sum output, the PEs in row 1 with the Upper and Lower outlier halves cannot compute the outlier partial sum output. After receiving outputs from row 1, ReCoN reorders and calculates the FP-outlier partial sum.
5
The controller signals PE row 2 to expect input partial sums from ReCoN and not from the previous row. If PE row 1 is the last row as in Figure 5(c), the oActs are directed to the oAct buffer.
6
oActs are post-processed (scaled and quantized). The post-processing of oActs does not require all the oActs to be computed. The post processing operations can be conducted as and when the oActs are generated and overlapped with the rest of the computation so as to hide computation and memory access latency overhead.
Figure 6. Microarchitecture of (a) Multi-precision PE, (b) Synchronization buffer and c) ReCoN switch.
@@ 5.2. On-chip Storage and Control
MicroScopiQ Controller. The controller generates appropriate control and signals to all units. It is functionally very similar to standard systolic array controllers (Sharma et al., 2018), however, with added functionality to support specific features of ReCoN (arbitration of simultaneous access by multiple rows), multi-precision PEs, post-processing unit. Furthermore, unlike traditional controllers, the MicroScopiQ controller exerts fine-grained control using handshaking signals on the streaming of iActs, iAccs into the PE array to account for the increased pipeline depth through ReCoN for rows with outliers.
5.3. Multi-Precision PE Array
Multi-Precision. We depict the multi-precision MicroScopiQ PE in Figure 6(a). Existing multi-precision accelerators (Guo et al., 2023; Sharma et al., 2018; Guo et al., 2022) typically follow a bottom-up approach, employing low-precision PEs and grouping neighboring PEs to support higher precision. This sacrifices throughput for multi-precision support, as multiple PE columns are required to perform a single MAC operation. This reduces parallelism and increases latency. We adopt a different strategy via the MODE signal for multi-precision support by mapping multiple weights that share the same rightbound iAct to the same PE at lower precision (2-bits) or a single weight at higher precision (4-bits). The two 2-bit weights mapped to the same PE in MicroScopiQ are the weights that would have been mapped to different columns within the same row in existing multi-precision accelerators. At high-precision operation we utilize available parallelism from the PE array while at lower-precision we increase throughput by the parallel evaluation of multiple partial sums.
Multiplication Stage. Inspired by (Liu et al., 2023; Umuroglu et al., 2018) we present a multiplier-tree architecture for multi-precision INT multiplication. The weights and iAct are partitioned across the four $4$-bit $\times$ $2$-bit multipliers to calculate partial sums ($P_{00},P_{01},P_{10},P_{11}$) as described in Figure 6(a). Based on the MODE signal the partial sums above are combined using a combination of adders and shifters to provide the result (Res) of different bit-precision weights with iAct as follows,